Electronic Journal of Qualitative Theory of Differential Equations 2009, No. 51, 1-1; http://www.math.u-szeged.hu/ejqtde/

Addendum to integrable and continuous solutions of a nonlinear quadratic integral equation

El-Sayed A.M.A and Hashem H.H.G Faculty of Science, Alexandria University, Alexandria, Egypt e.mails : amasayed@hotmail.com, amasayed5@yahoo.com

August 17, 2009

This addendum concerns the paper of the above title found in EJQTDE No. 25 (2008). There are some misprints in that paper:

(i) Page 3, line 5 should be k : [0, 1]×[0, 1] → R_{+} satisfies Carathèodory condition (i.e. measurable in t for all s ∈ [0, 1] and continuous in s for all t ∈ [0, 1]) such that ∫_{0}^{1} k(t, s)m_{2}(s)ds is bounded ∀t ∈ [0, 1].

(ii) Page 6, line 6 should be k : [0, 1]×[0, 1] → R_{+} satisfies Carathèodory condition (i.e. measurable in s for all t ∈ [0, 1] and continuous in t for all s ∈ [0, 1]) such that k(t, s)m_{2}(s) ∈ L_{1} ∀t ∈ [0, 1].

EJQTDE, 2009 No. 51, p. 1